1. One painter can paint an entire house in four hours. A second painter takes two hours on his own . If they paint the same house at the same time, how long would it take the two painters to paint one house?

Answer by nerdybill(7384)

1. One painter can paint an entire house in four hours. A second painter takes two hours on his own . If they paint the same house at the same time, how long would it take the two painters to paint one house?
.
Let t = time (hours) it takes for both
then
t (1/4 + 1/2) = 1
t (1 + 2) = 4
t (3) = 4
t = 4/3 hours
or
t = 1 hour and 20 minutes